OpenStudy (anonymous):

lim x-> pi/4 of (1-tanx)/(sinx-cosx)

OpenStudy (anonymous):

\[\lim_{x \rightarrow \pi/4}\] \[(1-\tan x)/(\sin x-\cos x)\]

OpenStudy (anonymous):

i'm not very good at limits but tan pi/4 = 1 so 1- tanx -- > 0

OpenStudy (jamesj):

\[ 1 - \tan x = 1 - (\sin x / \cos x) = \frac{\cos x - \sin x}{\cos x} \] Given that, can you see how to proceed?

OpenStudy (anonymous):

how did you get the cos x - sin x in the numerator?

OpenStudy (jamesj):

Cmon. I'm writing down the expression with a common denominator, cos x

OpenStudy (jamesj):

\[1 = \frac{ \cos x}{ \cos x} \]

OpenStudy (anonymous):

thanks. so is the answer - root2

OpenStudy (anonymous):

/

OpenStudy (jamesj):

exactly.

OpenStudy (anonymous):

yayy! thanks